What are the monthly payments (principal and interest) on a 30-year home mortgage for an $130,000 loan when interest rates are fixed at 6 percent? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

Monthly payments   
Answer #1
Monthly payments $ 779.42
Working:
# 1 Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.005)^-360)/0.005 i 6%/12 = 0.005
= 166.7916144 n 30*12 = 360
# 2 Monthly payment = Loan Amount / Present Value of annuity of 1
= $       1,30,000 / 166.7916
= $           779.42